### v1

@@ III.3 $\lambda$-trajectories for the pentaquark $P_{cc\bar{c}bb}$
The $\lambda_{1}$- and $\lambda_{2}$-trajectories for the pentaquark $P_{cc\bar{c}bb}$ in configurations $(cc)(\bar{c}(bb))$ and $(bb)(\bar{c}(cc))$ are shown in Fig. 3. The complete forms of the $\lambda$-Regge trajectories can be calculated easily from Eqs. (II.2) and (II.2) or Eqs. (II.2) and (II.2). Since the diquark $1$ and the triquark are treated as individual constituents without considering their internal substructures, the complete form of the $\lambda_{1}$-trajectories coincides with its main part. Thus, the $\lambda_{1}$-trajectories are the simplest among the four series of Regge trajectories (see Table 5 and Fig. 3). In Fig. 3, the graphs of the compete forms, the fitted formulas, and the main parts of the $\lambda_{1}$-trajectories overlap. Similar to the $\rho$-trajectories, the explicit forms of the $\lambda_{2}$-trajectories are lengthy. By performing a linear fit to the calculated data, we obtain the fitted formulas presented in Table 5. As shown in Fig. 3, these fitted formulas closely overlap with the complete forms of the $\lambda_{2}$-trajectories, indicating that the full forms can be well approximated by the simple fitted formulas. Therefore, both the complete forms and the fitted formulas for the $\lambda_{2}$-trajectories exhibit the behavior $M{\sim{x_{\lambda_{2}}^{2/3}}}$. However, the main parts of the complete forms of the $\lambda_{2}$-trajectories [listed in Table 5] deviate from the full expressions.
Table 4: Comparison of theoretical predictions for the spin-averaged masse of the ground state of pentaquark $P_{cc\bar{c}bb}$ (in GeV).
Mass
$\{cc\}(\bar{c}\{bb\})$ 14.13
$\{bb\}(\bar{c}\{cc\})$ 14.10
$\{bc\}(\bar{c}\{bc\})$ 14.13
Ref. Rashmi:2024ako 15.21
Ref. Liang:2024met 14.62
Ref. Gordillo:2024blx 14.30
Ref. An:2022fvs 14.57
Ref. Zhang:2023hmg 14.86
@@ IV Discussions
By employing the Regge trajectory formulas in Eqs. (II.2) and (II.2) or Eqs. (II.2) and (II.2), the masses of higher excited states can be easily calculated. We argue that calculating the masses of highly excited states is instructive, even though these states are expected to be unstable. This is because various models or approaches for calculating the masses of pentaquarks often employ different parameter values. Nevertheless, they generally provide similar predictions for the ground state and lower excited states. However, the theoretical predictions for the higher excited states are expected to differ more significantly, which can be instructive.
@@
It is interesting that the main parts of the complete forms and the fitted formulas have the same functional form $M=m_{R}+\beta(x+c_{0})^{2/3}$. However, the values of the parameters $m_{R}$ and $\beta$ change significantly. This means that, unlike in the meson case, an obvious and direct dependence of $m_{R}$ and $\beta$ on the constituents’ masses and string tension is no longer present when we construct simple $\rho_{1}$-, $\rho_{2}$-, and $\lambda_{2}$-trajectory formulas for pentaquarks. It can be concluded from the preceding discussions, as well as those in Sec. III, that the $\rho_{1}$, $\rho_{2}$, and $\lambda_{2}$ Regge trajectories for fully heavy pentaquarks cannot be obtained merely by mimicking the meson Regge trajectories; instead, they should be constructed based on the pentaquark’s structure and substructure. Otherwise, the $\rho_{1}$-, $\rho_{2}$-, and $\lambda_{2}$-trajectories must rely solely on fitting existing theoretical or future experimental data. The fundamental relationship between the slopes of the obtained trajectories and constituents’ masses and string tension will remain unclear. The predictive power of the Regge trajectories would be compromised.
@@
For fully heavy pentaquarks, all four series of Regge trajectories are concave downward in the $(M,\,x)$ planes when the confining potential is linear. It can also be easily shown that these trajectories are concave downward in the $(M^{2},\,x)$ planes when the confining potential is linear.
@@ V Conclusions
The complete forms of the Regge trajectories for pentaquark $P_{cc\bar{c}bb}$ are often lengthy and cumbersome. Except for the $\lambda_{1}$-trajectories, the $\rho_{1}$-, $\rho_{2}$-, and $\lambda_{2}$-trajectories for the fully heavy pentaquarks cannot be obtained by simply mimicking the meson Regge trajectories but should instead be constructed based on pentaquark’s structure and substructure. Otherwise, the $\rho_{1}$-, $\rho_{2}$-, and $\lambda_{2}$-trajectories must rely solely on fitting existing theoretical or future experimental data. The fundamental relationship between the slopes of the obtained trajectories and constituents’ masses and string tension will remain unclear. The predictive power of the Regge trajectories would be compromised.
@@ Appendix A List of the Regge trajectory relations
×$\times\left(0.334-\frac{0.97227\left(2.95+0.487766(0.205+n_{r_{1}})^{2/3}\right)}{14.1255+0.487766(0.205+n_{r_{1}})^{2/3}}\right)^{2/3}\left(1.008+\frac{0.0894041\left(2.95+0.487766(0.205+n_{r_{1}})^{2/3}\right)}{14.1255+0.487766(0.205+n_{r_{1}})^{2/3}}\right).$ (20)
@@
×$\times\left(1.008+\frac{0.0124(9.61+0.332799(0.01+n_{r_{2}})^{2/3})}{11.16+0.332799(0.01+n_{r_{2}})^{2/3}}\right)+0.486744\Big{(}\Big{(}14.1296+0.332799(0.01+n_{r_{2}})^{2/3}$

### v3

@@ III.3 $\lambda$-trajectories for the pentaquark $P_{cc\bar{c}bb}$
Fig. 3 shows the radial and orbital $\lambda_{1}$- and $\lambda_{2}$-trajectories for the pentaquark $P_{cc\bar{c}bb}$ in the configurations $(cc)(\bar{c}(bb))$ and $(bb)(\bar{c}(cc))$. Circles represent the predicted data, calculated using Eqs. (II.2) and (II.2) or Eqs. (II.2) and (II.2), as listed in Table 3. The black dashed lines represent the complete forms of the $\lambda$-trajectories, obtained from Eqs. (II.2) and (II.2) or from Eqs. (II.2) and (II.2). The pink lines represent the fitted formulas, obtained by a linear fit to the calculated data in Table 3 and listed in Table 6. The blue dotted lines correspond to the main parts of the full forms, which are also given in Table 6. Since the diquark $1$ and the triquark are treated as individual constituents without considering their internal substructures, the $\lambda_{1}$-trajectories are the simplest among the four series of Regge trajectories (see Table 6 and Fig. 3). In Fig. 3, the black dashed lines, pink lines, and blue dotted lines for the $\lambda_{1}$-trajectories overlap; that is, the compete forms, fitted formulas, and main parts are identical for these trajectories. Similar to the $\rho$-trajectories, the explicit forms of the $\lambda_{2}$-trajectories are lengthy. Performing a linear fit to the calculated data yields the fitted formulas in Table 6. As shown in Fig. 3, these fitted formulas closely match the complete forms of the $\lambda_{2}$-trajectories, indicating that the full forms can be well approximated by the simple fitted formulas. Therefore, both the complete forms and the fitted formulas for the $\lambda_{2}$-trajectories exhibit the behavior $M{\sim{x_{\lambda_{2}}^{2/3}}}$. However, the main parts for the complete forms of the $\lambda_{2}$-trajectories (listed in Table 6) deviate from the full expressions.
Table 4: Masses of the ground and first excited states of the pentaquark $P_{cc\bar{c}bb}$. The notation in Eq. (1) is expressed as $|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N_{2}^{2s_{3}+1}L_{2J_{2}},N_{1}^{2s_{4}+1}L_{1J_{1}}\rangle$. $|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N_{2}L_{2},N_{1}L_{1}\rangle$ denotes the spin-averaged states. The data are taken from Tables 2 and 3. $\dagger$} denotes that for the first five states, the angular momenta for different structures are all equal to zero, $l_{1}=l_{2}=L_{1}=L_{2}=0$. $\ddagger$} denotes that for the $|1^{3}s_{1},1^{3}s_{1},1P,1S\rangle$ state, $l_{1}=l_{2}=L_{1}=0$, $L_{2}=1$. $\ast$} denotes that for the $|1^{3}s_{1},1^{3}s_{1},1S,1P\rangle$ state, $l_{1}=l_{2}=L_{2}=0$, $L_{1}=1$.
$J^{P}$ $|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N_{2}L_{2},N_{1}L_{1}\rangle$ $(cc)(\bar{c}(bb))$ [GeV] $(bb)(\bar{c}(cc))$ [GeV]
$\left(\frac{1}{2}\right)^{-}$, $\left(\frac{3}{2}\right)^{-}$, $\left(\frac{5}{2}\right)^{-}$ $|1^{3}s_{1},1^{3}s_{1},1S,1S\rangle$ 14.13 14.10
$|2^{3}s_{1},1^{3}s_{1},1S,1S\rangle$ 14.49 14.41
$|1^{3}s_{1},2^{3}s_{1},1S,1S\rangle$ 14.45 14.46
$|1^{3}s_{1},1^{3}s_{1},2S,1S\rangle$ 14.45 14.43
$|1^{3}s_{1},1^{3}s_{1},1S,2S\rangle$ 14.58 14.54
$\left(\frac{1}{2}\right)^{+}$, $\left(\frac{3}{2}\right)^{+}$, $\left(\frac{5}{2}\right)^{+}$, $\left(\frac{7}{2}\right)^{+}$ $|1^{3}s_{1},1^{3}s_{1},1P,1S\rangle$ 14.36 14.33
$\left(\frac{1}{2}\right)^{+}$, $\left(\frac{3}{2}\right)^{+}$, $\left(\frac{5}{2}\right)^{+}$, $\left(\frac{7}{2}\right)^{+}$ $|1^{3}s_{1},1^{3}s_{1},1S,1P\rangle$ 14.45 14.42
For ease of comparison with possible experimental results and other theoretical approaches, the masses of the ground and first excited states of the pentaquark $P_{cc\bar{c}bb}$ are summarized in Table 4. For the first five states, the angular momenta for different structures are all zero. In general, the mass of the pentaquark $\left(\frac{1}{2}\right)^{-}$ state formed by a spin-$\frac{1}{2}$ triquark and a spin-1 diquark is not equal to that of the pentaquark $\left(\frac{1}{2}\right)^{-}$ state formed by a spin-$\frac{3}{2}$ triquark and a spin-1 diquark, although the mass difference is small. The same applies for the pentaquark $\left(\frac{3}{2}\right)^{-}$ state. For the $|1^{3}s_{1},1^{3}s_{1},1P,1S\rangle$ state, the pentaquark $\left(\frac{1}{2}\right)^{+}$ state can be formed by the spin-1 diquark and either a spin-$\frac{1}{2}$ triquark with $s_{3}=\frac{1}{2}$, a spin-$\frac{3}{2}$ triquark with $s_{3}=\frac{1}{2}$, a spin-$\frac{1}{2}$ triquark with $s_{3}=\frac{3}{2}$, or a spin-$\frac{3}{2}$ triquark with $s_{3}=\frac{3}{2}$. For the $|1^{3}s_{1},1^{3}s_{1},1S,1P\rangle$ state, the pentaquark $\left(\frac{1}{2}\right)^{+}$ state can be formed by the spin-1 diquark and either a spin-$\frac{1}{2}$ triquark with $s_{4}=\frac{1}{2}$, a spin-$\frac{1}{2}$ triquark with $s_{4}=\frac{3}{2}$, a spin-$\frac{3}{2}$ triquark with $s_{4}=\frac{1}{2}$, or a spin-$\frac{3}{2}$ triquark with $s_{4}=\frac{3}{2}$. The masses of these penatquark $\left(\frac{1}{2}\right)^{+}$ states will differ from one another; however, the mass differences among these states will be small. Other states can be discussed in a similar manner. These mass differences are relativistic corrections, which do not affect the Regge trajectory behavior. They are not considered here but will be addressed in future work.
@@ IV Discussions
To date, there is no evidence of a fully heavy pentaquark. However, many theoretical studies have been conducted on fully heavy pentaquarks Gordillo:2024blx ; Liang:2024met ; Gordillo:2023tnz ; An:2020jix ; An:2022fvs ; Sharma:2025adr ; Yan:2021glh ; Zhang:2023hmg ; Wang:2021xao ; Azizi:2025fmx ; Zhang:2020vpz ; Yang:2022bfu ; Rashmi:2024ako . These theoretical studies will aid in the experimental search for fully heavy pentaquarks. In this work, we present the $\rho$- and $\lambda$-trajectories for the fully heavy pentaquark $P_{cc\bar{c}bb}$, and apply the pentaquark Regge trajectory to provide a rough estimate of the masses of $P_{cc\bar{c}bb}$ (see Tables 2 and 3). The pentaquark Regge trajectory, given by Eqs. (II.2) and (II.2), or alternatively by Eqs. (II.2) and (II.2), can also be applied to fully charmed pentaquarks and other fully heavy pentaquarks. Note that when $\mu_{\lambda}>3.83$ GeV, the relations in Eq. (17) should be adjusted by fitting to experimental and theoretical data Xie:2024dfe .
Table 5: Comparison of theoretical predictions for the spin-averaged masse of the ground state of pentaquark $P_{cc\bar{c}bb}$.
Mass [GeV] Method
$\{cc\}(\bar{c}\{bb\})$ 14.13 Regge trajectory
$\{bb\}(\bar{c}\{cc\})$ 14.10 Regge trajectory
$\{bc\}(\bar{c}\{bc\})$ 14.13 Regge trajectory
Ref. Rashmi:2024ako 15.21 Effective quark mass
and screened charge scheme
Ref. Liang:2024met 14.62 Nonrelativistic potential quark model
Ref. Gordillo:2024blx 14.30 Schrödinger equation
Ref. An:2022fvs 14.57 Constituent quark model
Ref. Zhang:2023hmg 14.86 MIT bag model
@@
It is noteworthy that the main parts of the complete forms and the fitted formulas share the same functional form, $M=m_{R}+\beta(x+c_{0})^{2/3}$. However, the values of the parameters $m_{R}$ and $\beta$ differ considerably. For example, for the main part of the complete form of the $\rho_{2}$-trajectory, $m_{R}=13.8296$ and $\beta=0.332799$; these values are readily computed using Eqs. (II.2) and (II.2), and clearly depend on the constituents’ masses and the string tension. By contrast, for the fitted $\rho_{2}$-trajectory, $m_{R}=14.1131$ and $\beta=0.331141$; here, the dependence on the constituents’ masses and string tension is not readily apparent and becomes more complicated. This indicates that, unlike in the meson case, the dependence of $m_{R}$ and $\beta$ on the constituents’ masses and string tension is no longer obvious or direct when we construct simply fitted $\rho_{1}$-, $\rho_{2}$-, and $\lambda_{2}$-trajectory formulas for pentaquarks. As discussed above and in Sec. III, the $\rho_{1}$, $\rho_{2}$, and $\lambda_{2}$ Regge trajectories for fully heavy pentaquarks cannot be obtained by merely mimicking the meson Regge trajectories; instead, they should be constructed based on the actual structure and substructure of the pentaquark. Otherwise, the $\rho_{1}$-, $\rho_{2}$-, and $\lambda_{2}$-trajectories must rely solely on fitting existing theoretical or future experimental data. Consequently, the fundamental relationship between the slopes of the obtained trajectories and constituents’ masses and string tension will become unobvious and complicated. The predictive power of the Regge trajectories would be compromised.
@@
For potential models, both the dynamic equation and the potential determine the form of the Regge trajectories Chen:2022flh ; Chen:2021kfw ; Chen:2018bbr . Constituents are classified as light or heavy based on their masses, which correspond to ultrarelativistic or nonrelativistic dynamical equations, leading to different Regge trajectories Chen:2022flh ; Chen:2021kfw . At large $r$, the color Coulomb potential is neglected, while the confining potential plays a dominant role in potential (4), which affects the form of Regge trajectories brau:04bs ; Sonnenschein:2018fph . The parameters $\sigma$ and $C$ in Eq. (5) do not affect the forms of the Regge trajectories but do affect their values; for example, see Eqs. (II.2) and (II.2). For fully heavy pentaquarks, all four series of Regge trajectories are concave downward in the (M,x)(M,\,x) plane when the confining potential is linear. It can also be easily shown that these trajectories are concave downward in the $(M^{2},\,x)$ plane under the same conditions. For light systems, such as light mesons, the Cornell potential has the same functional form and, when the masses of the light constituents are neglected, results in linear Regge trajectories in the $(M^{2},\,x)$ plane Chen:2022flh ; Chen:2021kfw ; Lucha:1991vn . When the masses of the light constituents are taken into account, the Regge trajectories for light systems also become concave in the $(M^{2},\,x)$ plane Wilczek:2004im ; Chen:2023web . It is thus expected that the Regge trajectories for light, singly heavy, doubly heavy, triply heavy, quadruply heavy, and fully heavy pentaquarks are all concave downward in the $(M^{2},\,x)$ plane when the confining potential is linear and the masses of the light quarks are considered.
@@ V Conclusions
The complete forms of the Regge trajectories for pentaquark $P_{cc\bar{c}bb}$ are often lengthy and cumbersome. Except for the $\lambda_{1}$-trajectories, the $\rho_{1}$-, $\rho_{2}$-, and $\lambda_{2}$-trajectories for the fully heavy pentaquarks cannot be obtained by simply mimicking the meson Regge trajectories but should instead be constructed based on pentaquark’s structure and substructure. Otherwise, the $\rho_{1}$-, $\rho_{2}$-, and $\lambda_{2}$-trajectories must rely solely on fitting existing theoretical or future experimental data. The fundamental relationship between the slopes of the obtained trajectories and constituents’ masses and string tension will become unobvious. The predictive power of the Regge trajectories would be compromised.
@@
Acknowledgments
We are very grateful to the anonymous referees for the valuable comments and suggestions.
@@ Appendix A List of the Regge trajectory relations
$\times\left(0.334-\frac{0.97227\left(2.95+0.487766(0.205+n_{r_{1}})^{2/3}\right)}{14.1255+0.487766(0.205+n_{r_{1}})^{2/3}}\right)^{2/3}\left(1.008+\frac{0.0894041\left(2.95+0.487766(0.205+n_{r_{1}})^{2/3}\right)}{14.1255+0.487766(0.205+n_{r_{1}})^{2/3}}\right).$ (21)
@@
$\times\left(1.008+\frac{0.0124(9.61+0.332799(0.01+n_{r_{2}})^{2/3})}{11.16+0.332799(0.01+n_{r_{2}})^{2/3}}\right)+0.486744\Big(\Big(14.1296+0.332799(0.01+n_{r_{2}})^{2/3}$
@@
The $\rho_{1}$- and $\rho_{2}$-excited masses for the configuration $(cc)(\bar{c}(bb))$ are calculated using Eqs. (A) and (A), respectively, and are listed in Table 2. By linearly fitting the calculated masses in the $(M,(c_{0}+x)^{2/3})$ plane, we obtain the fitted formulas:
$M=13.9661+0.465794(0.205+n_{r_{1}})^{2/3}$ (23)
and
$M=14.1131+0.331141(0.01+n_{r_{2}})^{2/3},$ (24)
respectively. The complete forms of different Regge trajectories exhibit varying complexity [see Eqs. (A) and (A)]. However, their fitted counterparts, as shown in Eqs. (23) and (24), are comparatively simple. Moreover, $m_{R}$ and slope for the fitted Regge trajectories [Eqs. (23) and (24)] are different from those for the main parts of the complete forms of the Regge trajectories [see Eqs. (A) and (A) or Table 6.]. For the $\rho_{1}$-trajectory, $m_{R}=13.9661$ for the fitted form, while it is $13.8255$ for the main part. The values of $\beta_{n_{r_{1}}}$ are $0.465794$ for the fitted $\rho_{1}$-trajectory and $0.487766$ for the main part of the complete form, respectively.
@@
Data Availability Statement. This article has no associated data or the data will not be deposited.
Code Availability Statement. This article has no associated code or the code will not be deposited.
Open Access. This article is distributed under the terms of the Creative Commons Attribution License (CC-BY4.0), which permits any use, distribution and reproduction in any medium, provided the original author(s) and source are credited.